What is the complementary angle of (x+3)

Answer by Theo(13342)

two angles are complementary if their sum is 90 degrees.

if their sum is 90 degrees, then one of the angles is equal to 90 - the other angle.

for example:

30 + 60 = 90, therefore 30 and 60 are complementary.

30 = 90 - 60, therefore one of the angles (30) is equal to 90 - the other angle (60).

likewise 60 = 90 - 30.

in general:

if a + b = 90, then b = 90 - a and a = 90 - b.

your problem states to find the complementary angle of x + 3.

let y be that angle.

in order to be complmentary to x + 3, y would have to be equal to 90 - (x + 3)

y = 90 - (x + 3) becomes y = 90 - x - 3 which becomes y = 87 - x.

the complementary angle to x + 3 is equal to 87 - x.

if these angles are complementary, then their sum is 90 degrees.

x + 3 + 87 - x is equal to 90, therefore the angles are complementary.

your solution is that the angle that is complementary to x + 3 is 87 - x.